OpenStudy (anonymous):

The following pictures depict objects whose sizes have different orders of magnitude. The approximate sizes are indicated in meters, with negative powers of 10. The sizes in the above sequence make a geometric sequence. What is the ratio of the sequence? A. 0.01 B. 0.1 C. 1 D. 100 E. 10

OpenStudy (anonymous):

would this be 10

OpenStudy (anonymous):

i mean d

OpenStudy (butterflydreamer):

ratio = 2nd term / 1st term So the 2nd term = 10^-2 which is the same as 0.01 1st term = 1 So ratio = 0.01/1 = ?

OpenStudy (anonymous):

0.01 right

OpenStudy (butterflydreamer):

yeep

OpenStudy (anonymous):

wait how tho

OpenStudy (butterflydreamer):

look at your picture... what is the first term in your sequence? What is the second term in your sequence?

OpenStudy (anonymous):

1meter and 10^-2m

OpenStudy (butterflydreamer):

yeppp. To find the ratio you do 2nd term divided by 1st term... So 10^-2 / 1 = ?

OpenStudy (anonymous):

0.01

OpenStudy (butterflydreamer):

yep :) that's why we got 0.01

OpenStudy (anonymous):

ok i get it now thx